OpenStudy (anonymous):

multiply. (sqrt7)(-6sqrt7 + 4)

OpenStudy (anonymous):

-38??

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{7}\left(-6\sqrt{7}+4\right)\] \[\Large \sqrt{7}\left(-6\sqrt{7}\right)+\sqrt{7}\left(4\right)\] \[\Large -6\sqrt{7*7}+4\sqrt{7}\] \[\Large -6\sqrt{49}+4\sqrt{7}\] \[\Large -6\left(7\right)+4\sqrt{7}\] \[\Large -42+4\sqrt{7}\] So \[\Large \sqrt{7}\left(-6\sqrt{7}+4\right) = -42+4\sqrt{7}\]